# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v1

@@ Appendix D Success Event
$F_{t}^{\alpha}$ :=$:=\left\{\exists s,a,h,i:\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\geq\zeta_{h}^{t}(s,a)\right\},$
$F_{t}^{r}$ :=$:=\left\{\exists s,a,h:\left|\hat{r}_{h}^{t}(s,a)-r_{h}(s,a)\right|\geq\phi_{h}^{r,t}(s,a)\right\},$
$F_{t}^{g}$ :=$:=\left\{\exists s,a,h,i:\left|\hat{g}_{i,h}^{t}(s,a)-g_{i,h}(s,a)\right|\geq\phi_{i,h}^{g,t}(s,a)\right\},$
$F_{t}^{p}$ :=$:=\left\{\exists s,a,s^{\prime},h:\left|p_{h}(s^{\prime}\mid s,a)-\hat{p}_{h}^{t-1}(s^{\prime}\mid s,a)\right|\geq\phi_{h}^{p,t}(s,a,s^{\prime})\right\},$
$F_{t}^{N}$ :=$:=\left\{\exists s,a,h:N_{h}^{t}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right)\right\}.$
@@ Appendix E Policy Evaluation
5:         $\hat{Q}_{h}^{\pi_{t}}(s,a;\overline{g}_{i},\overline{p})=\min\Big{\{}\overline{g}_{i,h}(s,a)+\overline{p}_{h}(\cdot\mid s,a)\hat{V}_{\overline{g},h+1}^{\pi}(\cdot),H-h+1\Big{\}}$  ($\forall i\in[m]$)
@@ Appendix G Optimality and Optimism of Policies
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to time step $h$ in episode $t$. Suppose that for every $(t,h)$, $\zeta_{h}^{t}(s_{h},a_{h})$ satisfies
$\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})]\mid\mathcal{F}_{h}^{t-1}\right|\leq 2H.$
Then, with probability $1-\delta$,
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\left(\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid\mathcal{F}_{h}^{t-1}\right]\right)\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right).$
@@
for $t\in[T]$ and $h\in[H]$. For each $k\in[TH]$, set
$D_{k}:=\zeta_{k}(s_{h},a_{h})-\mathbb{E}[\zeta_{k}(s_{h},a_{h})\mid\mathcal{F}_{h}^{t-1}].$
which holds that $\mathcal{F}_{h}^{t}\subset\mathcal{F}_{h^{\prime}}^{t^{\prime}}$ for any $t\leq t^{\prime}$.
For any $k(t,h)\in[T]\times[H]$,
$\mathbb{E}[D_{k}\mid\mathcal{F}_{h}^{t-1}]=0,$
which means that $\{D_{k}\}_{k}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{h}^{t-1}\}$ and satisfies $|D_{k}|\leq 2H$. By the Azuma-Hoeffding inequality, for any $\epsilon>0$,
$\mathbb{P}\left(\sum_{k=1}^{T\cdot H}D_{k}>\epsilon\right)\leq 2\text{exp}\left(-\frac{\epsilon^{2}}{2TH\cdot(2H)^{2}}\right)=2\text{exp}\left(-\frac{\epsilon^{2}}{8TH^{3}}\right).$
Setting the right-hand side equal to $\delta$ yields $\epsilon=2\sqrt{2TH^{3}\text{ln}\frac{2}{\delta}}$. Then, with probability $1-\delta$, we have
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid\mathcal{F}_{h}^{t-1}\right]\right|=\left|\sum_{k=1}^{T\cdot H}D_{k}\right|\leq 2\sqrt{2TH^{3}\text{ln}\frac{2}{\delta}}\lesssim\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}.$
@@
$N_{h}^{(W_{t}),t}(s,a)>\frac{1}{2}\sum_{j=\max\{1,t-W_{t}+1\}}^{t-1}q_{h}^{\pi_{t}}(s,a;p)-H\ln\frac{SAH}{\delta^{\prime}},$
@@
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t}(s,a),1\}}}\mid\mathcal{F}_{t-1}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T/\gamma}+SAH).$
@@
We first define $w_{h}^{t}(s,a)=\mathbb{P}(s_{h}^{t}=s,a_{h}^{t}=a|\mathcal{F}_{t-1})$. Then, we have
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{{N_{h}^{(W_{t}),t}(s,a),1\}}}}\mid\mathcal{F}_{t-1}\right]$ $=\sum_{t,h,s,a}w_{h}^{t}(s,a)\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t}(s,a),1\}}}$
@@ G.1 Optimality under Pessimistic Thresholds
Lemma 3 (Policy optimality under pessimistic thresholds).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$ and parameter $\gamma\in(0,1]$, if we use pessimistic thresholds, then
@@
By applying Lemmas 14 and 15, and summing over $t$ and $h$, we further have
@@
$\lesssim m\rho\left[\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid s_{1},\pi,p)\right]+\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right]$
@@ G.2 Optimality of Optimistic Policies
Lemma 5 (Policy optimality with optimistic thresholds).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$, if we adopt the optimistic thresholds, then
@@ H.1 Proof of Theorem 2
Let $S$ be the number of states, $A$ be the number of actions, $H$ be the horizon length, $m$ be the number of constraints, $\rho$ be the Slater gap under pessimistic thresholds, $\gamma$ be the Growing-Window parameter, $T$ be the number of episodes and $\mathcal{N}$ denote the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability $1-\delta$, Algorithm 2 under pessimistic thresholds obtains the following regret bounds:
@@
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\frac{1}{\rho}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
$\leq$ $\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{T}\lambda+\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)+\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}.$
Under the pessimistic thresholds, we have $y_{t}\leq\tilde{y}_{t}$ at episode $t$, thus
$\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})\lambda\leq 0,$
leading us to simplify the above inequality as
$\mathcal{L}(\lambda)\leq\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)}_{(a)}+\underbrace{\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})}_{(b)}+\underbrace{\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}}_{(c)}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}.$ (13)
@@
Second term (reward estimation error)
@@
Third term (dual update bound)
@@
Substituting these results back into (13) and choosing the learning rate $\eta_{\lambda}=\sqrt{\frac{mH^{2}T}{\rho^{2}}}$, we achieve the intermediate bound, for any $0\leq\lambda\leq\rho\mathbf{1}$,
$\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{T}\lambda\lesssim$ $(\rho+\frac{\|\lambda\|_{2}^{2}}{\rho})\sqrt{mH^{2}T}+\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)$
@@
$\lesssim$ $\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}:=\theta(T).$ (15)
Due to the linearity of both the reward and constraints, combined with the convexity of the occupancy-measure set, we leverage Lemmas 24 and 25 and then derive the following result for any $T^{\prime}\in[T]$:
@@
$\mathcal{R}_{T}(g)\leq\tilde{\mathcal{O}}\left(\frac{1}{\rho}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right).$
@@ H.2 Proof of Theorem 3
Under the same conditions of Theorem 2, with probability $1-\delta$, Algorithm 2 under optimistic thresholds holds the following regret bounds:
@@
∎ By Lemmas 24 and 25, we have the following result for any $T^{\prime}\in[T]$: $\max_{i\in[m]}\left[\sum_{t=1}^{T^{\prime}}y_{t}\right]\leq\frac{\theta^{\prime}(T)}{\rho^{\prime}}$. Therefore, we conclude the bound for constraint violation

### v2

@@ Appendix D Success Event
$F_{t}^{\alpha}$ :=$:=\left\{\exists s,a,h,i:\left|\hat{\alpha}_{i,h}^{t-1}(s,a)-\alpha_{i,h}(s,a)\right|\geq\zeta_{h}^{t-1}(s,a)\right\},$
$F_{t}^{r}$ :=$:=\left\{\exists s,a,h:\left|\hat{r}_{h}^{t-1}(s,a)-r_{h}(s,a)\right|\geq\phi_{h}^{r,t-1}(s,a)\right\},$
$F_{t}^{g}$ :=$:=\left\{\exists s,a,h,i:\left|\hat{g}_{i,h}^{t-1}(s,a)-g_{i,h}(s,a)\right|\geq\phi_{i,h}^{g,t-1}(s,a)\right\},$
$F_{t}^{p}$ :=$:=\left\{\exists s,a,s^{\prime},h:\left\|p_{h}(s^{\prime}\!\mid\!s,a)-\hat{p}_{h}^{t-1}(s^{\prime}\!\mid\!s,a)\right\|_{1}H\geq\phi_{h}^{p,t-1}(s,a,s^{\prime})\right\},$
$F_{t}^{N}$ :=$:=\left\{\exists s,a,h:N_{h}^{t-1}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right)\right\}.$
@@ Appendix E Policy Evaluation
5:         $\hat{Q}_{h}^{\pi_{t}}(s,a;\overline{g}_{i},\overline{p})=\min\Big{\{}\overline{g}_{i,h}(s,a)+\overline{p}_{h}(\cdot\mid s,a)\hat{V}_{\overline{g},h+1}^{\pi_{t}}(\cdot),H-h+1\Big{\}}$  ($\forall i\in[m]$)
@@ Appendix G Optimality and Optimism of Policies
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to time step $h$ in episode $t$. Suppose that for every $(t,h)$, the term $\zeta_{h}^{t}(s_{h},a_{h})$ satisfies
$\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})\!\mid\!\mathcal{F}_{h-1}^{t}]\right|\leq 2H.$
Then, for any confidence parameter $\delta\in(0,1)$, with probability at least $1-\delta$,
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\left(\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\!\mid\!\mathcal{F}_{h-1}^{t}\right]\right)\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right).$
@@
for $t\in[T]$ and $h\in[H]$ and relabel the filtration by $\mathcal{F}_{k}=\mathcal{F}_{h}^{t}$ and $\mathcal{F}_{0}:=\{\emptyset,\Omega\}$. For each $k\in[TH]$, set
$D_{k}:=\zeta_{k}(s_{h},a_{h})-\mathbb{E}[\zeta_{k}(s_{h},a_{h})\mid\mathcal{F}_{k-1}].$
which holds that $\mathcal{F}_{k^{\prime}}\subset\mathcal{F}_{k}$ for any $k^{\prime}\leq k$.
For any $k\in[TH]$,
$\mathbb{E}[D_{k}\mid\mathcal{F}_{k-1}]=0,$
which means that $\{D_{k}\}_{k}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{k-1}\}$ and satisfies $|D_{k}|\leq 2H$. By the Azuma-Hoeffding inequality, for any $\epsilon>0$,
$\mathbb{P}\left(\left|\sum_{k=1}^{T\cdot H}D_{k}\right|>\epsilon\right)\leq 2\text{exp}\left(-\frac{\epsilon^{2}}{2TH\cdot(2H)^{2}}\right)=2\text{exp}\left(-\frac{\epsilon^{2}}{8TH^{3}}\right).$
Setting the right-hand side equal to $\delta$ yields $\epsilon=2\sqrt{2TH^{3}\text{ln}\frac{2}{\delta}}$. Then, with probability at least $1-\delta$, we have
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\left(\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid\mathcal{F}_{h-1}^{t}\right]\right)\right|=\left|\sum_{k=1}^{T\cdot H}D_{k}\right|\leq 2\sqrt{2TH^{3}\text{ln}\frac{2}{\delta}}\lesssim\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}.$
@@
$N_{h}^{(W_{t}),t}(s,a)>\frac{1}{2}\sum_{j=\max\{1,t-W_{t}+1\}}^{t}q_{h}^{\pi_{t}}(s,a;p)-H\ln\frac{SAH}{\delta^{\prime}},$
@@
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t}(s,a),1\}}}\mid\mathcal{F}_{t}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T/\gamma}+SAH).$
@@
We first define $w_{h}^{t}(s,a)=\mathbb{P}\{s_{h}^{t}=s,a_{h}^{t}=a\}$. Then, we have
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{{N_{h}^{(W_{t}),t}(s,a),1\}}}}\mid\mathcal{F}_{t}\right]$ $=\sum_{t,h,s,a}w_{h}^{t}(s,a)\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t}(s,a),1\}}}$
@@ G.1 Optimality under Pessimistic Thresholds
Lemma 3 (Policy optimality of pessimistic policies).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$ and parameter $\gamma\in(0,1]$, under pessimistic thresholds, it holds
@@
By applying Lemmas 4 and 14, and summing over $t$ and $h$, we further have
@@
$\lesssim m\rho\left[\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid s_{1},\pi,p\right]+\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right]$
@@ G.2 Optimality of Optimistic Policies
Lemma 5 (Policy optimality of optimistic policies).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$, under optimistic thresholds, the following holds:
@@ H.1 Proof of Theorem 2
Let $S$, $A$, $m$, $H$, $T$ be the number of states, actions, constraints, horizon length, and episodes, respectively. Define $\rho$ as the Slater gap under pessimistic thresholds, $\gamma$ as the Growing-Window parameter, and $\mathcal{N}$ as the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability at least $1-\delta$, Algorithm 2 under pessimistic thresholds achieves the following regret bounds:
@@
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\left(1+\frac{1}{\rho}\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
$\leq$ $\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)}_{(a)}+\underbrace{\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda}_{(b)}+\underbrace{\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})}_{(c)}+\underbrace{\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}}_{(d)}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}.$ (14)
@@
Second term (constraint and threshold estimation error)
In the pessimistic case, it holds that $\alpha_{i}\leq\overline{\alpha}_{t,i}$ for any constraint $i$ and episode $t$. We then obtain
$\left|\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right|$ $=\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)+\sum_{t=1}^{T^{\prime}}\left(\alpha_{i}-\overline{\alpha}_{t,i}\right)\right|$
$\leq\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|.$
According to Equation (H.2), we have
$(b)=\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda\leq$ $\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\|\lambda\|_{2}$
$\leq$ $\sqrt{m}\|\lambda\|_{2}\max_{i\in[m]}\left|\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right|$
$\leq$ $\sqrt{m}\|\lambda\|_{2}\max_{i\in[m]}\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|$
$\leq$ $\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)\|\lambda\|_{2}.$
Third term (reward estimation error)
@@
Fourth term (dual update bound)
@@
Substituting these results back into (H.1) and choosing the learning rate $\eta_{\lambda}=\sqrt{\frac{mH^{2}T}{\rho^{2}}}$, we achieve the intermediate bound, for any $0\leq\lambda\leq\rho\mathbf{1}$,
$\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{\top}\lambda\lesssim$ $(\rho+\frac{\|\lambda\|_{2}^{2}}{\rho})\sqrt{mH^{2}T}+\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)$
$+\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)\|\lambda\|_{2}$
@@
$\lesssim$ $\left(1+\rho\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}:=\theta(T).$ (16)
Due to the linearity of both the reward and constraints, combined with the convexity of the occupancy-measure set, we leverage Lemmas 23 and 24 and then derive the following result for any $T^{\prime}\in[T]$:
@@
$\mathcal{R}_{T}(g)\leq\tilde{\mathcal{O}}\left(\left(1+\frac{1}{\rho}\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right).$
@@ H.2 Proof of Theorem 3
Under the same conditions as Theorem 2, with probability at least $1-\delta$, Algorithm 2 under optimistic thresholds guarantees the following regret bounds:
@@
where $\rho^{\prime}$ is the Slater gap under optimistic thresholds.
@@
By Lemmas 23 and 24, we have the following result for any $T^{\prime}\in[T]$: $\max_{i\in[m]}\left[\sum_{t=1}^{T^{\prime}}y_{t}\right]\leq\frac{\theta^{\prime}(T)}{\rho^{\prime}}$. Therefore, we conclude the bound for constraint violation